# Predicting the Geolocation of Tweets Using transformer models on Customized Data

### v3

@@ Abstract
This research is aimed to solve the tweet/user geolocation prediction task and provide a flexible methodology for the geotagging of textual big data. The suggested approach implements neural networks for natural language processing (NLP) to estimate the location as coordinate pairs (longitude, latitude) and two-dimensional Gaussian Mixture Models (GMMs). The scope of proposed models has been finetuned on a Twitter dataset using pretrained Bidirectional Encoder Representations from Transformers (BERT) as base models. Performance metrics show a median error of fewer than 30 km on a worldwide-level, and fewer than 15 km on the US-level datasets for the models trained and evaluated on text features of tweets’ content and metadata context. Our source code and data are available at https://github.com/K4TEL/geo-twitter.git
@@ 1 Introduction
Twitter, being a widely used online social network, accumulates a large volume of diverse data at a high velocity, this includes short and disordered tweets, a vast network of users and rich contextual information for both users and tweets. This data serves as input for studying three common geolocation problems such as user home location, tweet location and mentioned location prediction [54]. Although, much research dedicated to predicting users’ home location utilizes the user network of connections (e. g. followings, replies, mentions) to achieve higher accuracy [56] [55] [53], this work focuses primarily on textual data analysis. Therefore, only tweet content (the text itself) and tweet context (tweet and user meta-data) are adopted to solve the problem of tweet location prediction.
@@
The solving of the tweet location prediction problem requires the labeled location which is commonly extracted from the tweet’s meta-data. In terms of Twitter, only 1-2% of the tweets are geotagged with longitude-latitude coordinates [34] which is by default followed by the accurate place description (e. g. country, city, place name). Therefore, there are two types of location definition - a pair of numerical coordinates and a textual class label - that imply the type of a task to be either a numerical regression or a textual label classification. The latter is commonly applied for relatively small-scale problems such as location prediction on a city level [22]. However, this study focuses on the country and global worldwide-level scale problems, thus the former method using longitude-latitude coordinates as a labeled location was chosen as more appropriate. Since Twitter geolocation data is stored in a standard format of a coordinate system called the World Geodetic System 1984 (WGS84), the same geodetic datum was utilized to represent the tweets’ labeled location. Although the WGS84 is a continuous, global reference system, when data is displayed on a two-dimensional map, such as a Mercator projection, numerical values of coordinates stay in strict ranges:
$Y=(y_{lon},y_{lat});\quad y_{lon}\in[-180,180];\quad y_{lat}\in[-90,90]$
The decision to solve the numerical regression rather than the label classification task using BERT-based models was also supported by the experiment results presented in Scherrer’s work [42] showing that regression outperformed classification in the geospatial accuracy (mean and median error distance) metrics for the limited Twitter datasets of BosnianCroatian-Montenegrin-Serbian and Deutch languages.
@@ 2 Related works
The most straightforward approach is to analyze the natural languages present in social media posts. There are multiple projects processing documents or the summary of user profile tweets to predict the author geolocation [47] [38] [48] [18] [31]. These works apply geographical grids to divide the Earth’s surface into different-sized regions such that highly populated areas are split into smaller pieces while more sparse population areas end up as large grid cells. Early efforts [15] [38] were mainly focused on mining indicative information from users’ posting content relying on location indicative words (LIWs) that can link users to their home locations, based on various NLP techniques (e.g., topic models and statistic models) [56]. For example, Rahimi extracts bag-of-words features from user posts [36]; Wing and Baldridge estimate the word distributions for different regions [47]. Other word-centric works [39] [5] [25] [43] [33] also focus on filtering LIWs from the text and using the gazetteers including not only city/country names but also dialect terms to resolve the geographical indexes. While such methodologies suppose the predefined set of LIWs and their geographical coordinates, Rahimi [35] proposes a neural network encoding of such words and phrases to the continuous two-dimensional space. In the oldest of relevant works, Eisenstein presents similar unsupervised models which are topic-based [12] [13]. In this light, we aimed to develop another unsupervised method based on the state-of-the-art models for NLP with no need for LIWs datasets or geographic gazetteers.
@@
Overall, studies in the previous works show that the prediction of user home locations draws equal significance from both Twitter network structures and tweet content [37] [11] [56] [55] [53]. However, it’s noteworthy that only a limited number of studies [4] [40] [52] leverage Twitter network data for predicting tweet locations.
@@
Additionally, users can enrich their profiles with details like hometowns, timezones, personal websites, and bios. However, common methods mainly focus on content information due to the deficient data records obtained from publicly available sources of Twitter geotagged datasets. Since the usage of context is not well explored for tweet location prediction [54], there was an opportunity for data feature engineering of tweets’ metadata contained in our raw datasets of archived Twitter posts.
Some researchers also define metadata features as model inputs, usually in addition to network and textual inputs, in order to predict users’ home locations [10] [32] [17]. Among utilized metadata varieties, temporal information like tweet timestamps and user-declared timezones are shown as effective in implying tweet and home locations at coarse-grained granularity [7] [14] [24]. For instance, Yuan in the work [51] managed to cover the temporal aspect and users’ mobility to estimate each tweet’s location source. However, the inputs of numerical type, such as timestamps and timezone are more ambiguous for the text-oriented transformer models explored in this study, therefore, time-related features were mainly ignored on account of more relevant input options. It’s important to note that the user’s timezone metadata is no longer used in Twitter, but would only provide insight into the geographic region of the user’s home country or its capital city.
@@
In the realm of user home location prediction, some authors came up with probabilistic models solving the classification task on the country/city levels [16] [27] [28]. However, employing such classification-based approaches becomes impractical for tweet location prediction, given the potential presence of hundreds of thousands of Points of Interest (POIs) within a city. While classification-based methodologies are commonly employed in home location prediction, the same does not hold true for coordinate-oriented tweet location prediction. Thus, there are relatively few studies that adopt a classification-based approach to predict the location of individual tweets in the form of specific POIs, as exemplified by [4] and [18].
Importantly, Priedhorsky’s methodology was aligned with the present work as it utilized GMMs for geolocation prediction instead of simple coordinates prediction [34]. Another subsequent study, which employed a hybrid approach, also estimated geolocation through GMMs rather than coordinates [2]. This study used the text and network features jointly as predictors for the final estimation. The performance metrics used in both works were adopted to evaluate the proposed probabilistic models as described in Section C.
@@
The base BERT model is intended to be fine-tuned on a downstream task which, in this case, was a regression to geographical coordinates (and GMM parameters such as weights and covariance) output as a type of sentence classification task. In general, BERT models have been pretrained using Masked Language Modeling (MLM), which enables the model to learn bidirectional representations of sentences, and Next Sentence Prediction (NSP) to capture the relationships between sentences. Since BERT’s layers are hierarchical, early BERT layers learn more generic linguistic patterns, such as differences between multiple languages and their general sentence structures (in the case of the BERT multilingual model). While the later layers learn more task-specific patterns, in this case - the associations between points on the world map and specific terms referred to as LIWs, as well as text constructions and linguistic patterns commonly used in certain geographical areas.
@@
Such constraints posed a challenge in accurately classifying a user’s home location, as concatenating all of the user’s tweets into a single text would result in an input exceeding the limit. As a result, users’ location prediction was performed on a per-tweet basis, and it was observed that the majority vote approach frequently (42%) led to misclassification of the user’s home location. Thus only the task of tweet geolocation prediction could be efficiently solved using the BERT-base models. In contrast, this work focuses on processing smaller (less than 300 words) text samples and aggregating a set of probabilistic predictions in a form of GMMs to estimate the set of most probable user’s home location points as described in Section LABEL:sec:user-predict.
@@
In this work, the geographic granularity level was defined as a worldwide area covering the whole scope of countries and languages present in the Twitter feed. However, the results of Scherrer’s study [42] demonstrated that language-specific BERT models clearly outperforms their multilingual counterpart on most of the used datasets. In light of these findings, the country-level model for the United States was finetuned on the US dataset utilizing the primarily English BERT base model, technically referred to as "bert-base-cased". Nevertheless, the global dataset necessitated the use of the multilingual BERT base model ("bert-base-multilingual-cased"), which resulted in higher error distance metrics for the per-tweet and per-user geolocation prediction, as expected. A critical challenge addressed is the representation of multiple languages, since the ability to process multiple languages is a key factor in dividing text inputs into geospatial regions. To tackle this issue, a multilingual BERT model, pretrained on the 104 largest Wikipedia languages, was utilized as a starting point for finetuning on a global dataset with geolocation labels.
@@ 3 Materials and Methods
Considering that the two-dimensional distributions were used as the output format of the probabilistic models, this work proposes the custom procedures of loss function computation for the four model types determined in Table 3 based on the model output format. The training procedure for the best of the proposed models is demonstrated on the flowchart diagram in Figure 1. In terms of the input data used for model fine-tuning, this study put into service only textual information such as the tweet content (raw user’s text) and tweet context (metadata like information from user’s profile and geo-tag textual descriptions).
@@
Step 1 - parameterization of the model architecture, model outputs, and finetuning; Step 2 - importing default tokenizer and base BERT model; Step 3 - reading dataset files; Step 4 - forming text input features; Step 5 - tokenization of the text inputs; Step 6 - forming of the model data loader; Step 7 - loading input features to the model; Step 8 - text processing by base model; Step 9 - returning raw model outputs; Step 10 - comparing processed predictions and labels; Step 11 - computed loss back-propagation.
The regression layer was used to convert the vector of final hidden states made up of 768 floats to the specified number of numerical outputs. According to the task of geolocation prediction, the essential number of outputs was defined as 2 standing for the geographic point (longitude, latitude) in the WGS84 coordinate system. However, the best of the proposed model had an output of 20 numerical values utilized as parameters for the definition of predicted GMM. Note that the models predicting a single coordinate pair could be trained using the standard Mean Squared Error (MSE) loss function. However, more complex output forms such as GMMs required a custom loss function described in Section 3.2. The Probabilistic Single/Multiple Outcome Prediction (PSOP/PMOP) models have slightly outperformed the straightforward approach of Geospatial Single/Multiple Outcome Prediction (GSOP/GMOP) models on the median error distance metrics as shown in Table 4.
The hyper-parameters tuning stage has covered such parameters as the type of the scheduler, minimal and maximal learning rates for the chosen scheduler, and the number of epochs. For the current task, the cosine type was chosen from the set of linear, cyclic, step, and plateau schedulers. Experiments have shown that the optimal learning rate reduction range starts at 1e-5 and ends at 1e-6 at the end of the last training epoch. The entering number of epochs was reduced from 5 to 3 for the dataset consisting of 2.7 million training samples, which are grouped into batches of 10 to 16 tweets in the model data loader. This decision was made as a result of the observed lack of significant error distance reduction after the third epoch for both training and test metrics. By default, the development set of 300,000 samples was evaluated without a gradient propagation at the end of each epoch. The described parameters were held constant across all proposed models, with variations only occurring in the loss function type, the number of outcomes (prediction points), covariance type, and the combination of text features.
@@ 3.1 Data preprocessing
The datasets used for finetuning the models consisted of tweets collected between the years 2020-2022 which have defined "coordinates" and "place" objects in the tweet JSON. The datasets have undergone several preprocessing stages prior to being uploaded into the model, including text strings filtering, rearrangement of the columns to form text features, tokenization, and split into tensor data loaders of preset batch size. The model was operating solely with input IDs and attention masks representing text features of the dataset and its numerical labels which corresponded to geolocation coordinates.
Generally, only 5% of all tweets in our Twitter archive database collected from 2020 to 2022 have geolocation coordinates, such that the number of available geotagged samples was approximately 150 million. It has been found that up to 20% of the geotagged tweets in the collected datasets were posted by bot users, which are automated accounts e.g., newspapers, weather forecasts, airplane trackers, etc. While the finetuning datasets contained tweets of both real and bot users, the evaluation datasets were filtered out based on the assumption that real users don’t post more than 20 messages per day.
The training data we collected for this study includes a comprehensive set of tweets, their metadata (author profiles, location geotags), and labeled location coordinates serving as the input features and "ground truth" for our model’s learning process. During the preprocessing stage, the original tweet objects were condensed into the text content ("text"), relevant metadata context ("user" and "place"), and geolocation coordinates ("lon", "lat") representing the labeled location.
@@
Fortunately, Twitter metadata stored in the "place" field of geo-tagged tweets provides automatically generated information about the country, country code, place type, location name, and location full name. Since tweet JSON root-level objects “place” and “user” had multiple fields containing potentially relevant context data, some of them were used in forming the input features as shown in Table 1. Despite the fact that only 26% of users provide location information in their profiles according to [6] and that user-generated data can be noisy, the proposed multitask learning procedures enhanced the geospatial accuracy and resulted in the median error of fewer than 30km.
@@
Since typical model input shouldn’t include "place" context which refers exclusively to geo-tagged tweets, multitask learning has been implemented in order to both preserve the input format and utilize "place" metadata during training. This approach implies separate wrapper layers for each of the textual input features and a per-batch loss common for all features during finetuning. The main input feature for the best of the proposed models included the tweet text Content combined with context obtained from the author profile (username, description, and location), while the supplemental input feature was dedicated solely to the Context data obtained from the "place" field.
Considering that evaluation was performed on the typical model input ("text" and "user"), models trained with the described Key and Minor Feature setup showed lower spatial errors in comparison to the models trained without the minor feature or the models trained on a combination of all data in a single text feature ("text", "user", and "place"). Tweet metadata provided context for the Content of the tweet and has been shown to improve overall model accuracy, as demonstrated in Table 4 by the underperformance of the "TO" (TEXT-ONLY) model trained only on the "text" contents without metadata.
@@
The labeled location for this study was defined as the pair of longitude-latitude coordinates associated with each geotagged tweet. By definition, there are two JSON children entities that describe the location information of a tweet: "coordinates" and "place". The former is present only when the exact location was assigned by the user, therefore data labels in a longitude-latitude format have been collected by parsing only the "coordinates" object of tweets.
In contrast to the tweet numerical geotag, the "place" field provides an accurate automatically generated by Twitter textual description of the location, such as country-city naming, which was utilized as a supplementary input during the finetuning of the best models. As shown in Table 1, the fields of the "place" object utilized in forming the training dataset included country, place type, name, location, and full name of the geotagged location.
The decision to split the training input into Key and Minor Features was dictated by the assumption about the model’s attention shift. The model trained on the combination of all input features was expected to learn to pay attention only to the geotags. To avoid memorizing of positions of the geotag tokens in the input sequence during training, the randomization of words could not be applied due to the semantics distortion.
The evaluation results of "A" (ALL) and "NG+GO" (NON-GEO) models shown in Table 4 demonstrate the difference between the model trained on all input features combined into one and the same model of 5 prediction points trained on the split data (NON-GEO and GEO-ONLY). The latter clearly outperformed the former, thus supporting the idea of applying multitask learning techniques involving multiple input text features.
There was a limited number of text feature combinations explored in this study as described in Table 2. Note that each training sample had separate tokenized sequences for all input features assigned to it in the model data loader, such that per-batch loss of a training step depended on each feature prediction component as shown in Figure 9.
@@
Table 2: Text features combinations used in text-based (Content) and hybrid (Content and Context) approaches
The models were evaluated on diverse arbitrary tweets imitating a common situation when the "place" field is unavailable and only the key features of always present “text” and “user” inputs are used for prediction. In general, each tweet always contains “text” and “user” (NON-GEO) used as the Key training Feature input during the evaluation of test and validation datasets.
In practice, each model had a necessary Key Feature (KF) and an optional Minor Feature (MF) wrapper layers which were used according to the input type during the finetuning stage. Utilizing the "place" field as MF and user-generated text ("text" and "user") as KF was crucial for predicting geolocation coordinates accurately. The "place" field provided valuable complementary information, enhancing context understanding and disambiguating similar locations, while tweet content offered relevant cues and linguistic patterns, improving prediction precision.
The described combination of NON-GEO KF and GEO-ONLY MF created a robust and flexible model, capable of handling diverse tweets when the "place" field is unreliable or unavailable. Therefore, the resulting trained system possesses the capability to effectively analyze and predict the labeled location of tweets based on any arbitrary input.

### v4

@@ Abstract
This research is aimed to solve the tweet/user geolocation prediction task and provide a flexible methodology for the geo-tagging of textual big data. The suggested approach implements neural networks for natural language processing (NLP) to estimate the location as coordinate pairs (longitude, latitude) and two-dimensional Gaussian Mixture Models (GMMs). The scope of proposed models has been finetuned on a Twitter dataset using pretrained Bidirectional Encoder Representations from Transformers (BERT) as base models. Performance metrics show a median error of fewer than 30 km on a worldwide-level, and fewer than 15 km on the US-level datasets for the models trained and evaluated on text features of tweets’ content and metadata context. Our source code and data are available at github.com/geo-twitter
@@ 1 Introduction
Twitter, being a widely used online social network, accumulates a large volume of diverse data at a high velocity, this includes short and disordered tweets, a vast network of users and rich contextual information for both users and tweets. This data serves as input for studying three common geolocation problems such as user home location, tweet location and mentioned location prediction [44]. Although, much research dedicated to predicting users’ home location utilizes the user network of connections (e. g. followings, replies, mentions) to achieve higher accuracy [46] [45] [43], this work focuses primarily on textual data analysis. Therefore, only tweet content (the text itself) and tweet context (tweet and user metadata) are adopted to solve the problem of tweet location prediction.
@@
The solving of the tweet location prediction problem requires the labeled location coordinates which are commonly extracted from the tweet’s metadata. In terms of Twitter, only 1–2% of the tweets are geo-tagged with longitude-latitude coordinates [24] which is by default followed by the accurate place description (e. g. country, city, place name). Therefore, there are two types of location definition: a pair of numerical coordinates and a textual class label—that imply the type of a task to be either a numerical regression or a textual label classification. The latter is commonly applied for relatively small-scale problems such as location prediction on a city level [16].
However, this study focuses on the country and global worldwide-level scale problems, thus the former method using longitude-latitude coordinates as a labeled location was chosen as more appropriate. Since Twitter geolocation data is stored in a standard format of a coordinate system called the World Geodetic System 1984 (WGS84), the same geodetic datum was utilized to represent the tweets’ labeled location. Although the WGS84 is a continuous, global reference system, when data is displayed on a two-dimensional map, such as a simple Plate carrée projection, where numerical values of longitude and latitude stay in strict ranges of 360 and 180 degrees respectively.
The decision to solve the numerical regression rather than the label classification task using BERT-based models was also supported by the experiment results presented in Scherrer’s work [31] showing that regression outperformed classification in the geospatial accuracy (mean and median error distance similar to SAE described in Section B.1) metrics for the limited Twitter datasets of BosnianCroatian-Montenegrin-Serbian and Deutch languages. Detailed descriptions of the proposed regression wrapper layers, outputs post-processing, and loss computation procedures can be found in Section 3.
@@ 2 Related works
The most straightforward approach is to analyze the natural languages present in social media posts. There are multiple projects processing documents or the summary of user profile tweets to predict the author geolocation [28] [37] [11]. These works apply geographical grids to divide the Earth’s surface into different-sized regions such that highly populated areas are split into smaller pieces while more sparse population areas end up as large grid cells. Early efforts were mainly focused on mining indicative information from users’ posting content relying on location indicative words (LIWs) that can link users to their home locations, based on various NLP techniques (e.g., topic models and statistic models) [46]. For example, Rahimi extracts bag-of-words features from user posts [26]; Wing and Baldridge estimate the word distributions for different regions [36]. Other word-centric works [32] [23] [42] also focus on filtering LIWs from the text and using the gazetteers including not only city/country names but also dialect terms to resolve the geographical indexes. While such methodologies suppose the predefined set of LIWs and their geographical coordinates, Rahimi [25] proposes a neural network encoding of such words and phrases to the continuous two-dimensional space. In this light, we aimed to develop another semi-supervised method based on the state-of-the-art models for NLP with no need for LIWs datasets or geographic gazetteers.
@@
Overall, studies in the previous works show that the prediction of user home locations draws equal significance from both Twitter network structures and tweet content [27] [46] [45] [43]. However, it’s noteworthy that only a limited number of studies [4] [29] [41] leverage the Twitter network of user connections for predicting tweet locations.
@@
Additionally, users can enrich their profiles with details like hometowns, timezones, personal websites, and bios. However, common methods mainly focus on content information due to the deficient data records obtained from publicly available sources of Twitter geo-tagged datasets. Since the usage of context is not well explored for tweet location prediction [44], there was an opportunity for data feature engineering of tweets’ metadata contained in our raw datasets of archived Twitter posts.
Some researchers also define metadata features as model inputs, usually in addition to network and textual inputs, in order to predict users’ home locations [22] [10]. Among utilized metadata varieties, temporal information like tweet timestamps and user-declared timezones are shown as effective in implying tweet and home locations at coarse-grained granularity [8]. For instance, Yuan in the work [40] managed to cover the temporal aspect and users’ mobility to estimate each tweet’s location source. However, the inputs of numerical type, such as timestamps and timezone are more ambiguous for the text-oriented transformer models explored in this study, therefore, time-related features were mainly ignored on account of more relevant input options. It’s important to note that the user’s timezone metadata is no longer used in Twitter, but would only provide insight into the geographic region of the user’s home country or its capital city.
@@
In the realm of user home location prediction, some authors came up with probabilistic models solving the classification task on the country/city levels [9] [19]. However, employing such classification-based approaches becomes impractical for tweet location prediction, given the potential presence of hundreds of thousands of Points of Interest (POIs) within a city. While classification-based methodologies are commonly employed in home location prediction, the same does not hold true for coordinate-oriented tweet location prediction. Thus, there are relatively few studies that adopt a classification-based approach to predict the location of individual tweets in the form of specific POIs, as exemplified by [4] and [11].
Importantly, Priedhorsky’s methodology was aligned with the present work as it utilized GMMs for geolocation prediction instead of simple coordinates prediction [24]. Another subsequent study, which employed a hybrid approach, also estimated geolocation through GMMs rather than coordinates [2]. This study used the text and network features jointly as predictors for the final estimation. The performance metrics used in both works were adopted to evaluate the proposed probabilistic models as described in Section B.2.
@@
The base BERT model is intended to be finetuned on a downstream task which, in this case, was a regression to geographical coordinates (and GMM parameters such as weights and covariance) output as a type of sentence classification task. In general, BERT models have been pretrained using Masked Language Modeling (MLM), which enables the model to learn bidirectional representations of sentences, and Next Sentence Prediction (NSP) to capture the relationships between sentences. Since BERT’s layers are hierarchical, early BERT layers learn more generic linguistic patterns, such as differences between multiple languages and their general sentence structures (in the case of the BERT multilingual model). While the later layers learn more task-specific patterns, in this case—the associations between points on the world map and specific terms referred to as LIWs, as well as text constructions and linguistic patterns commonly used in certain geographical areas.
@@
Such constraints posed a challenge in accurately classifying a user’s home location, as concatenating all of the user’s tweets into a single text would result in an input exceeding the limit. As a result, users’ location prediction was performed on a per-tweet basis, and it was observed that the majority vote approach frequently (42%) led to misclassification of the user’s home location. Thus only the task of tweet geolocation prediction could be efficiently solved using the BERT-base models. In contrast, this work focuses on processing smaller (less than 300 words) text samples and aggregating a set of probabilistic predictions in a form of GMMs to estimate the set of most probable user’s home location points as described in Section A.
@@
In this work, the geographic granularity level was defined as a worldwide area covering the whole scope of countries and languages present in the Twitter feed. However, the results of Scherrer’s study [31] demonstrated that language-specific BERT models clearly outperforms their multilingual counterpart on most of the used datasets. In light of these findings, the country-level model for the English and French-speaking countries was finetuned on a limited subset of French, Canadian, and British tweets as described in Section 3.3.2.
Nevertheless, the global dataset necessitated the use of the multilingual BERT base model (’bert-base-multilingual-cased’), which resulted in higher error distance metrics for the per-tweet and per-user geolocation prediction, as expected. A critical challenge addressed is the representation of multiple languages, since the ability to process multiple languages is a key factor in dividing text inputs into geospatial regions. To tackle this issue, a multilingual BERT model, pretrained on the 104 largest Wikipedia languages, was utilized as a starting point for finetuning on a global dataset with geolocation labels.
@@ 3 Materials and Methods
Considering that the two-dimensional distributions were used as the output format of the probabilistic models, this work proposes the custom procedures of loss function computation for the four model types different in output formats. The training procedure for the best of the proposed models is demonstrated on the flowchart diagram in Figure 1. In terms of the input data used for model fine-tuning, this study put into service only textual information such as the tweet content (raw user’s text) and tweet context (metadata like information from user’s profile and geo-tag textual descriptions).
@@
Step 1: parameterization of the model’s architecture, outputs, and finetuning;
Step 2: importing default tokenizer and base BERT [7] model;
Step 3: reading dataset files from the Twitter archive into a virtual sheet;
Step 4: forming inputs NON-GEO as KF and GEO-ONLY as MF according to Section 3.1;
Step 5: tokenization of the text inputs through the standard BERT tokenizer;
Step 6: forming of the model data loaders by selection of batch subsets from data;
Step 7: loading input features to the model, each to separate wrapper layer;
Step 8: text processing by the local base BERT model before applying wrapper layers;
Step 9: returning defined in size model outputs according to Table 3;
Step 10: comparing processed predictions and labels according to Section C;
Step 11: backpropagation of the total per-batch loss to the local base BERT version.
The regression layer was used to convert the vector of final hidden states made up of 768 floats to the specified number of numerical outputs. According to the task of geolocation prediction, the essential number of outputs was defined as 2 standing for the geographic point (longitude, latitude) in the WGS84 coordinate system. However, the best of the proposed model had an output of 20 numerical values utilized as parameters for the definition of predicted GMM. Note that the models predicting a single coordinate pair could be trained using the standard Mean Squared Error (MSE) loss function. However, more complex output forms such as GMMs required a custom loss function described in Section 3.2. The intention was to test whether the Probabilistic Single/Multiple Outcome Prediction (PSOP/PMOP) models outperform the straightforward approach of Geospatial Single/Multiple Outcome Prediction (GSOP/GMOP) models.
The hyper-parameters tuning stage has covered such parameters as the type of the scheduler, minimal and maximal learning rates for the chosen scheduler, and the number of epochs. For the current task, the cosine type was chosen from the set of linear, cyclic, step, and plateau schedulers. Experiments have shown that the optimal learning rate reduction range starts at 1e-5 and ends at 1e-6 at the end of the last training epoch. The entering number of epochs was reduced from 5 to 3 for the dataset consisting of 2.7 million training samples, which are grouped into batches of 10 to 16 tweets in the model data loader. This decision was made as a result of the observed lack of significant error distance reduction after the third epoch for both training and test (development set) metrics. By default, the development set of 300,000 samples was evaluated without a gradient propagation at the end of each epoch. The described parameters were held constant across all proposed models, with variations only occurring in the loss function type, the number of outcomes (prediction points), covariance type, and the combination of text features.
@@ 3.1 Data preprocessing
The datasets used for finetuning the models consisted of tweets collected between the years 2020–2022 which have defined ’coordinates’ and ’place’ objects in the tweet JSON. The datasets have undergone several preprocessing stages prior to being uploaded into the model, including text strings filtering, rearrangement of the columns to form text features, tokenization, and split into tensor data loaders of preset batch size. The model was operating solely with input IDs and attention masks representing text features of the dataset and its numerical labels which corresponded to geolocation coordinates.
Generally, only 5% of all tweets in our Twitter archive database collected from 2020 to 2022 have geolocation coordinates, such that the number of available geo-tagged samples was approximately 150 million. It has been found that up to 20% of the geo-tagged tweets in the collected datasets were posted by bot users, which are automated accounts e.g., newspapers, weather forecasts, airplane trackers, etc. While the finetuning datasets contained tweets of both real and bot users, the evaluation datasets were filtered out based on the assumption that real users don’t post more than 20 messages per day.
The training data we collected for this study includes a comprehensive set of tweets and their metadata (author profiles, location geo-tags) serving as inputs, and labeled location coordinates playing role of the models’s ’ground truth’ through the training and evaluation. During the preprocessing stage, the original tweet objects were condensed into the text content (’text’), relevant metadata context (’user’ and ’place’), and geolocation coordinates (’lon’, ’lat’) representing the labeled location.
@@
Fortunately, Twitter metadata stored in the ’place’ field of geo-tagged tweets provides automatically generated information about the country, country code, place type, location name, and location full name. Since tweet JSON root-level objects “place” and “user” had multiple fields containing potentially relevant context data, some of them were used in forming the input features as shown in Table 1. Despite the fact that only 26% of users provide location information in their profiles according to [5] and that user-generated data can be noisy, the proposed multitask learning procedures ought to enhance the geospatial accuracy.
@@
Since typical model input shouldn’t include ’place’ context which refers exclusively to geo-tagged tweets, multitask learning has been implemented in order to both preserve the input format and utilize ’place’ metadata during training. This approach implies separate wrapper layers for each of the textual input features and a per-batch loss common for all features during finetuning. The main input feature for the best of the proposed models included the tweet text Content combined with context obtained from the author profile (username, description, and location), while the supplemental input feature was dedicated solely to the Context data obtained from the ’place’ field.
Considering that evaluation ought to be performed on the typical model input (’text’ and ’user’), models trained with the described Key and Minor Feature setup were expected to show lower spatial errors in comparison to the models trained without the minor feature or the models trained on a combination of all data in a single text feature (’text’, ’user’, and ’place’). Moreover, the tweet metadata providing context for the Content of the tweet was expected to improve overall model accuracy by comparison with the ’TO’ (TEXT-ONLY) model trained only on the ’text’ contents without metadata. More details on the comparison of merged and split input features can be found in Section 4 and specifically in Table 4 containing all performance measurements.
@@
The labeled location for this study was defined as the pair of longitude-latitude coordinates associated with each geo-tagged tweet. By definition, there are two JSON children entities that describe the location information of a tweet: ’coordinates’ and ’place’. The former is present only when the exact location was assigned by the user, therefore data labels in a longitude-latitude format have been collected by parsing only the ’coordinates’ object of tweets.
In contrast to the tweet numerical geo-tag, the ’place’ field provides an accurate automatically generated by Twitter textual description of the location, such as country-city naming, which was utilized as a supplementary input during the finetuning of the best models. As shown in Table 1, the fields of the ’place’ object utilized in forming the training dataset included country, place type, name, location, and full name of the geo-tagged location.
The decision to split the training input into Key and Minor Features was dictated by the assumption about the model’s attention shift. The model trained on the combination of all input features was expected to learn to pay attention only to the geo-tags. To avoid memorizing of positions of the geo-tag tokens in the input sequence during training, the randomization of words could not be applied due to the semantics distortion.
The comparison of ’A’ (ALL) and ’NG+GO’ (NON-GEO+GEO-ONLY) models was intended to prove that the identical model setups trained on merged and split input features would differ in accuracy. In case the separation of GEO-ONLY from the always available NON-GEO enhanced the NON-GEO accuracy, it would support the idea of applying multitask learning techniques in this work despite the slower in total processing of multiple inputs and outputs.
There was a limited number of text feature combinations explored in this study as described in Table 2. Note that each training sample had separate tokenized sequences for all input features assigned to it in the model data loader, such that per-batch loss was computed as a linear combination of per-feature errors. The visualization of the multitask loss computation procedure can be found in Figure 9 alongside the scheme for a single feature (Key only) model of K outcomes in Figure 8.
@@
Table 2: Text features combinations used in text-based (Content) and hybrid (Content and Context) approaches.
The models were evaluated on diverse arbitrary tweets imitating a common situation when the ’place’ field is unavailable and only the key features of always present “text” and “user” inputs are used for prediction. In general, each tweet always contains “text” and “user” (NON-GEO) used as the Key training Feature, and as an input during the evaluation of development and validation datasets.
In practice, each model had a necessary Key Feature (KF) and an optional Minor Feature (MF) wrapper layers which were used according to the input type during the finetuning stage. Utilizing the ’place’ field as MF and user-generated text (’text’ and ’user’) as KF was crucial for predicting geolocation coordinates accurately. The ’place’ field provided valuable complementary information, enhancing context understanding and disambiguating similar locations, while tweet content offered relevant cues and linguistic patterns, improving prediction precision.
The described combination of NON-GEO KF and GEO-ONLY MF created a robust and flexible model, capable of handling diverse tweets when the ’place’ field is unreliable or unavailable. Therefore, the resulting trained system possesses the capability to effectively analyze and predict the labeled location coordinates based on any arbitrary input of the tweet-like length.